OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6 a)1 b)4 c)5 d)25

OpenStudy (anonymous):

3x+3 =-2x+2+6 3x+2x = 2+6-3 proceed

OpenStudy (anonymous):

get a like terms together. like this 3x+2x=2+6-3 add the x terms together and number terms and the number terms together like this 5x=5 divide by 5 to get five 5 to get x on it is own , X would equal 1